According to the Air Transport Association of America, the average operating cost of an MD-80 jet airliner is $2,087 per hour. Suppose the operating costs of an MD-80 jet airliner are normally distributed with a standard deviation of $169 per hour. (Round the value of z to 2 decimal places. Round your answers to 2 decimal places.)

(a) At what operating cost would only 23% of the operating costs be less? $

(b) At what operating cost would 65% of the operating costs be more? $

(c) What operating cost would be more than 85% of operating costs?
Answer #1

1.
=Mean+Z*Standard deviation
=2087-0.74*169
=1961.94

2.
=Mean+Z*Standard deviation
=2087-0.39*169
=2021.09

3.
=Mean+Z*Standard deviation
=2087+1.04*169
=2262.76